OpenStudy (anonymous):

HELP!! Samantha puts an initial $500 into a savings account. The account has a 4% annual compound interest rate. What is the function that best represents Samantha's account and how much does she have in savings after 5 years? Round your answer to the nearest dollar. Hint: Use the formula, f(x) = P(1 + r)x. A)f(x)=500(1.04)^x, $2,689 B)f(x)=500(0.04)^x, $512 C)f(x)=500(1.04)^x, $608 D)f(x)=500(1.04)^x, $527

OpenStudy (anonymous):

This just seems like a plug and chug type of question. Taking the formula and information given already, what can you make from it?

OpenStudy (anonymous):

Well it is 500 * 1.04^5

OpenStudy (anonymous):

Which is 608$

OpenStudy (anonymous):

yeah, so it's as easy as putting that into a calculator

OpenStudy (jdoe0001):

hint: \(\bf \qquad \textit{Compound Interest Earned Amount}\\ A=f(x)=P\left(1+r\right)^{x} \qquad \begin{cases} A=\textit{current amount}\\ P=\textit{original amount deposited}\to &\$500\\ r=rate\to 4\%\to \frac{4}{100}\to &0.04\\ x=years\to &5 \end{cases}\)

OpenStudy (anonymous):

well he just gave you your answer of C) so...XD

OpenStudy (anonymous):

Yeah...XD thank you all for your help